# Giant Heat Flux Effect in Non-Chiral Transmission Lines

### v2

@@ Abstract
We develop a theory of heat transport in non-chiral transmission lines (TLs) of quantum Hall edge channels coupled to Ohmic contacts (OCs), where heat transport is driven by charge fluctuations in the heat Coulomb blockade regime. This framework challenges conventional paradigms by revealing a giant heat flux effect—a significant amplification in heat transport arising from non-trivial fluctuation-dissipation relations. Through a Langevin-based approach, we derive the effective noise power in the chiral currents, which underlies this enhanced heat flux. Our findings predict clear experimental signatures unique to non-chiral TLs as well as providing insights into finite-frequency effects, showing crossovers to more conventional diffusive behavior. This work offers a perspective on feedback mechanisms in quasi-1D heat transport with implications for dissipation control in low-dimensional quantum systems.
@@
An essential component of mesoscopic QH circuits is the Ohmic contact (OC), a floating metallic granule coupled to 1D ballistic QH edge channels at various filling factors [10, 11, 12]. It serves as a probe for electrical and thermal properties [13, 12, 14, 15, 16, 17], while also introducing Coulomb interactions between the edge states, characterized by an energy scale determined by the granule’s charging energy $E_{C}$ [18, 19, 20, 21]. At low temperatures where $T/E_{C}\ll 1$, the Coulomb barrier inhibits charge and heat transport in classical systems [22] and acts as a band-pass filter in quantum systems. This leads to exciting phenomena such as dynamical renormalization of charge [23, 24] and heat transport [25, 26], electron state teleportation [27, 28], Luttinger liquid behavior [29, 30], tunable multi-channel Kondo effects [31, 32], charge fractionalization [33, 34], and Coulomb-mediated heat drag [35].
@@
A series of Ohmic contacts (OCs) configured in a transmission line (TL) effectively modeled by floating, interacting reservoirs [36, 37, 38, 39], can represent disorder in QH edges due to QH puddle formation [40, 41]. This setup enables theoretical studies that yield experimentally verifiable predictions, including models explaining how equilibrium is reached in disordered QH edge states [42], demonstrating that energy transport remains quantized despite dissipation, and offering insights into more exotic phenomena such as “negative heat drag” [36, 43].
@@
Originally proposed to resolve the “missing heat” paradox [44, 5, 36, 43], this TL framework focuses on the strongly interacting regime where the OC charging energy is significant  [^1] Similar models have been suggested for fractional QH edges in the non-interacting limit [39, 37, 38], based on the Wiedemann-Franz law, which we demonstrate does not always apply.. Unexpectedly, it revealed unconventional physics, including a set of “magic” Lorenz numbers [46], edge states that carry excess heat beyond the quantum limit [47], and behaviors closely linked to fractional statistics, resembling anyonic excitations [34]. In our previous work [47], we demonstrated that introducing a feedback loop breaks local chirality, enabling charge fluctuations to create correlated states through a backaction effect.
In this letter, we investigate the natural extension of this phenomenon by examining globally non-chiral systems, specifically a non-chiral TL composed of a series of OCs, as illustrated in Fig. 1. We analyze this setup under four boundary conditions: open, where the TL connects to large macroscopic reservoirs without charge fluctuations at one or both ends; closed, where the TL is terminated at both ends; and periodic, forming a continuous loop. Refer to Fig. 1(b-e) for illustrations. A key distinction from other diffusive systems is our ability to individually probe the chiral current components in the TL, for example, through capacitive coupling or an attached quantum point contact.
@@
Figure 1: Schematic of a non-chiral transmission line (TL) comprising $N=3$ floating ohmic contacts (OCs) (green) connected to chiral one-dimensional quantum Hall edge channels (red arrows) at the boundary of a quantum Hall bulk with filling factor $\nu=1$ (light blue). The edge channels can extend from large metallic contacts (yellow) located at the left and right ends of the TL. The TL can be configured with various boundary conditions (b-e), including both open and closed system boundaries. Fluctuations in the edge channels induce charge fluctuations in the OC located at position $n$ in the TL, denoted as $\Delta Q_{n}$. These fluctuations generate charge current fluctuations $\delta I^{\text{L/R}}_{n}$ in the outgoing channels to the left and right of each OC. For closed system boundary conditions (d and e), there is an additional constraint that total charge is conserved, meaning $\sum_{n}\Delta Q_{n}(t)=0$ at all times. For illustration, panel (a) shows $N=3$ OCs in the case of a fully open setup (b).
@@
Our key results are: (i) a nontrivial fluctuation-dissipation relation (FDR) in which noise power is proportional to the current response function of a specific current source, revealed by individual probing of chiral currents; (ii) a giant heat flux in open systems that greatly exceeds the equilibrium quantum of heat flux $J_{q}=\frac{\pi}{12}T^{2}$; (iii) a reduced heat flux in closed systems, indicating impedance in heat transport; and (iv) nontrivial electrical and thermal tunneling conductances that modify the Lorenz number, providing a distinctive signature for observing these altered edge states.
@@
In this letter, we focus on the heat Coulomb blockade (HCB) regime, where the charging energy $E_{C}$ significantly exceeds the OC temperature $T\ll E_{C}$, or equivalently, $\tau_{C}T\ll 1$. By definition, fluctuations vanish on average: $\langle\delta I^{\sigma,\text{c}}_{n}(t)\rangle=0$ and $\langle\Delta Q_{n}(t)\rangle=0$. Inserting the Langevin equation (2) into Kirchhoff’s law (1) leads to a second-order inhomogeneous difference equation, which can be solved using a discrete Green’s function [48, 49, 50]:
@@
where the noise power $S^{\sigma}_{n}$ is given by Eq. (10), and $\lim_{T\rightarrow 0}S^{\sigma}_{n}(\omega)$ represents the subtraction of vacuum fluctuations. In the heat Coulomb blockade regime, characterized by $\tau_{C}T\ll 1$, the heat flux can be evaluated as:
@@
Our analysis so far assumes both the heat blockade regime, $\tau_{C}T\ll 1$, and the condition $\tau_{\text{th}}T\ll 1$, which imposes a stringent requirement, but allows us to neglect the frequency dependence of the response function. Beyond this regime, the finite frequency behavior shows a universal, boundary condition-independent temperature dependence $J^{\sigma}_{N/2}/J_{q}=3\zeta\left(\frac{3}{2}\right)/(\sqrt{2}\pi^{3/2})\left(\tau_{C}T\right)^{-1/2}$, summarized in Fig. 2. Analogous arguments apply to the other boundary conditions listed in Table 1.
@@
$L^{\sigma}_{n}(N)/L_{0}$ $\frac{3\left(N+1+2Nn-2n^{2}\right)}{3-2n^{2}+N(3+2n)}$ $\frac{3(2n+1)}{2n+3}$ $\frac{3\left(1+6n^{2}-3N-6nN+2N^{2}\right)}{1+6n^{2}+9N-6nN+2N^{2}}$ $\frac{3\left(N^{2}+5\right)}{N(N+12)+5}$
@@
This work opens several avenues: (i) extending the model to fractional QH edge states, which could probe strongly interacting, disordered systems with diverse edge reconstructions; (ii) deriving non-equilibrium properties and examining modifications to Fourier’s law; (iii) applying this framework to 2D TL lattices to study bulk and edge effects in disordered 2D metals; (iv) exploring 2D random networks where Ohmic contacts may alter percolation physics and localization; and (v) examining TLs with finite-transparency QPCs, which would illuminate transport properties at partial edge channel transmission.
F.S thanks M. Straub for fruitful discussions. The authors acknowledges the financial support from the Swiss National Science Foundation.
@@
for the auto and cross noise power $(n\neq m)$, respectively. The equilibrium noise powers are linear in the Green’s function, as expected for a Fluctuation-Dissipation relation.
@@ I Green’s Function at Finite Frequencies (Open System Boundary Conditions)
$J_{n}^{\sigma}-J_{q}=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\text{Re}\left[\frac{(r^{n}-r^{-n})(r^{N+1-m}-r^{-(N+1-m)})}{(r-r^{-1})(r^{N+1}-r^{-(N+1)})}\right]\left(S^{\sigma}_{n}(\omega)-\omega\theta(\omega)\right),$ (63)

### v3

@@ Abstract
We develop a theory of heat transport in non-chiral transmission lines (TLs) of quantum Hall edge channels coupled to Ohmic contacts (OCs) that accounts for a dynamical accumulation of charge in the reservoirs. As a consequence, heat transport is driven by charge fluctuations in the heat Coulomb blockade regime. This framework challenges conventional paradigms by revealing a giant heat flux effect—a significant amplification in heat transport arising from non-trivial fluctuation-dissipation relations. Through a Langevin-based approach, we derive the effective noise power in the chiral currents, which underlies this enhanced heat flux. Our findings predict clear experimental signatures unique to non-chiral TLs, as well as provide insights into finite-frequency effects, showing crossovers to more conventional diffusive behavior. This work offers a perspective on feedback mechanisms in quasi-1D heat transport with implications for dissipation control in low-dimensional quantum systems.
@@
An essential component of mesoscopic QH circuits is the Ohmic contact (OC), a floating metallic granule coupled to 1D ballistic QH edge channels at various filling factors [10, 11, 12]. It serves as a probe for electrical and thermal properties [13, 12, 14, 15, 16, 17], while also introducing Coulomb interactions between the edge states, characterized by an energy scale determined by the granule’s charging energy $E_{C}$ [18, 19, 20, 21]. At low temperatures where $T/E_{C}\ll 1$, the Coulomb barrier inhibits charge and heat transport in classical systems [22] and acts as a band-pass filter in quantum systems. This leads to exciting phenomena such as dynamical renormalization of charge [23, 24] and heat transport [25, 26], electron state teleportation [27, 28], Luttinger liquid behavior [29, 30], tunable multi-channel Kondo effects [31, 32], charge fractionalization [33, 34, 35, 36], and Coulomb-mediated heat drag [37].
@@
A series of chirally connected Ohmic contacts (OCs) configured in a transmission line (TL) effectively modeled by floating, interacting reservoirs [38], can be an effective model of disorder in QH edges due to QH puddle formation [39, 40]. This setup enables theoretical studies that yield experimentally verifiable predictions, including models explaining how equilibrium is reached in disordered edge states [41], demonstrating that energy transport remains quantized despite dissipation, and offering insights into more exotic phenomena such as “negative heat drag” [38, 42].
@@
Originally proposed to resolve the “missing heat” paradox [43, 5, 38, 42], the chiral TL framework focuses on the interacting regime where the OC charging energy is significant. Unexpectedly, it revealed unconventional physics, including a set of “magic” Lorenz numbers [44], edge states that carry excess heat beyond the quantum limit [45], and behaviors closely linked to fractional statistics, resembling anyonic excitations [34]. In [45], we demonstrated that introducing a feedback loop breaks local chirality, enabling charge fluctuations to create correlated states through a backaction effect, causing all of these exotic phenomena.
In this Letter, our motivation is two-fold: First, we extend the study of heat transport to globally non-chiral systems, a natural progression from previous work focused on chiral quantum Hall edge channels, expecting similar exotic physics. Specifically, we investigate a non-chiral TL composed of a series of OCs, as shown in Fig. 1, under four boundary conditions: open, closed, periodic, and hybrid configurations (Fig. 1(b-e)). Second, we contrast our approach with existing studies on non-chiral transmission lines, predominantly in the context of fractional QH states [46, 47, 48, 49, 50, 51]. Unlike prior work, we explicitly incorporate the dynamical accumulation of charges $Q_{n}(t)$ in reservoirs ($\partial_{t}Q_{n}(t)\neq 0$)—a critical aspect neglected in most models, which typically focus on zero-frequency correlation functions. However, we stress that in general, these effects cannot be neglected, as their relevance depends on the probed energy scales $T/E_{C}$ and $\tau_{\text{th}}T$, where the Thouless time $\tau_{\text{th}}$ will be discussed below.
Our setup offers another key advantage over other diffusive systems: the ability to individually probe the chiral current components in the TL. This can be achieved, for instance, through capacitive coupling or an attached quantum point contact, enabling precise exploration of fluctuation and response mechanisms.
@@
Figure 1: Schematic of a non-chiral transmission line (TL) comprising $N=3$ floating ohmic contacts (OCs) (green) connected to chiral one-dimensional quantum Hall edge channels (red arrows) at the boundary of a quantum Hall bulk with filling factor $\nu=1$ (light blue). The edge channels can extend from large metallic contacts (yellow) located at the left and right ends of the TL. The TL can be configured with various boundary conditions (b-e), including both open and closed system boundaries. Fluctuations in the edge channels induce charge fluctuations in the OC located at position $n$ in the TL, denoted as $\Delta Q_{n}$. These fluctuations generate charge current fluctuations $\delta I^{\text{L/R}}_{n}$ in the outgoing channels to the left and right of each OC. For closed system boundary conditions (d and e), there is an additional constraint that total charge is conserved, meaning $\sum_{n}\Delta Q_{n}(t)=0$ at all times. For illustration, panel (a) shows $N=3$ OCs in the case of a fully open setup (b). The orange dot schematically indicates a measurement of the chiral current component, here $\delta I_{n+1}^{R}$, using an adjacent quantum dot (QD) or quantum point contact (QPC).
@@
Our key results are: (i) a nontrivial fluctuation-dissipation relation (FDR) in which noise power is proportional to the current response function of a specific current source, revealed by individual probing of chiral currents; (ii) a giant heat flux in open systems that greatly exceeds the equilibrium quantum of heat flux $J_{q}=\frac{\pi}{12}T^{2}$; and (iii) nontrivial electrical and thermal tunneling conductances that modify the Lorenz number, providing a distinctive signature for observing these altered edge states.
@@
In this section, we focus on the heat Coulomb blockade (HCB) regime, where the charging energy $E_{C}$ significantly exceeds the OC temperature $T\ll E_{C}$, or equivalently, $\tau_{C}T\ll 1$. By definition, fluctuations vanish on average: $\langle\delta I^{\sigma,\text{c}}_{n}(t)\rangle=0$ and $\langle\Delta Q_{n}(t)\rangle=0$. Inserting the Langevin equation (2) into Kirchhoff’s law (1) leads to a second-order inhomogeneous difference equation, which can be solved using a discrete Green’s function [52, 53, 54]:
@@
where the noise power $S^{\sigma\sigma}_{nn}$ is given by Eq. (10), and $\lim_{T\rightarrow 0}S^{\sigma\sigma}_{nn}(\omega)$ represents the subtraction of vacuum fluctuations. In the heat Coulomb blockade regime, characterized by $\tau_{C}T\ll 1$, the heat flux can be evaluated as:
@@
Our analysis so far assumes both the heat blockade regime, $\tau_{C}T\ll 1$, and the condition $\tau_{\text{th}}T\ll 1$, which imposes a stringent requirement, but allows us to neglect the frequency dependence of the response function. Beyond this regime, the finite frequency behavior shows a universal, boundary condition-independent temperature dependence $\lim_{N\rightarrow\infty}J^{\sigma}_{N/2}/J_{q}=3\zeta\left(\frac{3}{2}\right)/(\sqrt{2}\pi^{3/2})\left(\tau_{C}T\right)^{-1/2}$, summarized in Fig. 2. Analogous arguments apply to the other boundary conditions listed in Table 1.
@@
$L^{\sigma}_{n}(N)/L_{0}$ $\frac{3\left(N+1+2Nn-2n^{2}\right)}{3-2n^{2}+N(3+2n)}$ $\frac{3(2n+\sigma)}{2+2n+\sigma}$ $\frac{3\left(1+6n^{2}-3N-6nN+2N^{2}\right)}{1+6n^{2}+9N-6nN+2N^{2}}$ $\frac{3\left(N^{2}+5\right)}{N(N+12)+5}$
@@
This work opens several avenues: (i) extending the model to fractional QH edge states, which could probe strongly interacting, disordered systems with diverse edge reconstructions; (ii) deriving non-equilibrium properties and examining modifications to Fourier’s law; (iii) applying this framework to 2D TL lattices to study bulk and edge effects in disordered 2D metals; (iv) exploring 2D random networks where Ohmic contacts may alter percolation physics and localization; and (v) examining TLs with finite-transparency QPCs, which would reveal more intricate Kondo physics compared to two site models currently investigated.
F.S. thanks M. Straub for fruitful discussions. The authors also thank C. Spånslätt, C. Glattli, and D. Karki for their useful comments on the manuscript. The authors acknowledges the financial support from the Swiss National Science Foundation.
@@
Supplemental Material for “Giant Heat Flux Effect in Non-Chiral Transmission Lines” Florian Stäbler Alioune Gadiaga Eugene V. Sukhorukov November 27, 2024
@@
for the auto and cross noise power $(n\neq m)$, respectively. The equilibrium noise powers are linear in the Green’s function, as expected for a Fluctuation-Dissipation relation. Note that the unsymmetrized noise power (17) is not necessarily real; however, when computing observables such as the heat flux, only the symmetric and thus real part of the noise power contributes.
@@ I Green’s Function at Finite Frequencies (Open System Boundary Conditions)
$J_{n}^{\sigma}-J_{q}=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\,\text{Re}\left[\frac{(r^{n}-r^{-n})(r^{N-n}-r^{-(N-n)})}{(r-r^{-1})(r^{N+1}-r^{-(N+1)})}\right]\left(S^{\sigma}_{n}(\omega)-\omega\theta(\omega)\right),$ (63)